If a $25 investment is worth $50 today, how long must that $25 have been invested at 6% interest compounded twice a year?
Answer by stanbon(75887)

If a $25 investment is worth $50 today, how long must that $25 have been invested at 6% interest compounded twice a year?
=====
50 = 25(1+(0.06/2))^(2t)
2 = (1.03)^2t
2t = log(2)/log(1.03)
2t = 23.45
t = 11.72 years
=====================
Cheers,
Stan H.